# Correlations for aerodynamic force coefficients of non-spherical particles in compressible flows

### v1

@@ Abstract
This study presents particle-resolved direct numerical simulations using three-dimensional body-fitted hexahedral grids to investigate the aerodynamic force and torque coefficients of non-spherical particles in compressible flows. The simulations focus on three particle shapes: a prolate spheroid, an oblate spheroid, and a rod-like particle, across a range of Mach numbers (0.3 to 2.0), angles of attack (0°to 90°), and particle Reynolds numbers (100 to 300). Results show that particle shape significantly impacts the aerodynamic forces on a particle in a compressible flow, with oblate spheroids exhibiting the highest drag, lift, and torque values. Correlations for these aerodynamic coefficients of the particles in a compressible flow are developed and validated. These correlations advance multiphase flow modeling by improving the accuracy of point-particle simulations for non-spherical particles in compressible flows.
@@ 1 Introduction
Fully-resolved direct numerical simulations (PR-DNS) of particle-laden flows require tremendous computational resources. However, a point-particle approach can significantly reduce the computational power required to model particle-laden flows. This approach requires force models to determine the forces exerted on the particles by the surrounding fluid, as well as by contact with other particles [26]. In this context, several force models have been developed which are applicable for various flow regimes and volume fractions [44, 50, 23, 24, 39, 38, 11, 17, 34, 9]. In order to derive these correlations, it is essential to first understand the behaviour of an isolated particle and the forces experienced by the particle in various flow regimes. These force correlations can then be extended to multi-particle systems.
@@
Forces acting on an isolated particle depend on several factors, such as the particle Reynolds number, the Mach number, the fluid properties, the particle shape, and angle of attack between the mean fluid flow and the axis of the particle. Predicting the forces that act on a particle in a fluid flow and predicting its behaviour have been the subject of studies for decades [43, 48, 17, 36, 19, 25, 44, 29, 30, 27, 47]. One of the first analytical models to predict the drag of a spherical particle for particle Reynolds numbers in the Stokes regime ($\mathrm{Re}_{\mathrm{p}}\ll 1$) was given by Stokes [43] as $C_{\mathrm{D}}=24/\mathrm{Re_{p}}$ where $C_{\mathrm{D}}$ is the drag coefficient. This model is the basis for later proposed drag models [22, 8, 16, 37], applicable to various flow regimes and practical applications. Since particles are not limited to spherical shapes, the proposed models also included non-spherical particle shapes, for instance using a sphericity parameter proposed by Wadell [48]. Although this has been a major development in predicting drag for various particle shapes, it has an obvious drawback: the sphericity factor does not take into account the shape of the particle and only determines how closely the shape of the particle resembles that of a perfect sphere. This means that the information on the angle of attack between the mean flow direction and the major axis of the particle and morphology in different directions is excluded from the derived correlations, thereby affecting the accuracy of the correlations. Subsequently, Rosendahl [36] introduced a new method to describe the shape of non-spherical particles. This multi-parameter approach, includes the parameters of the minor axis, the aspect ratio and the super-elliptic exponent, and allows the prediction of aerodynamic forces of particles with a more complex shape. Similarly, Loth [23] uses shape correlation parameters based on the aspect ratio of the particle to determine the drag coefficient. In addition to these studies, several other correlations for lift, drag, and torque have been proposed based on experimental and numerical data for various flow regimes and particle shapes. Some of the most notable correlations were those presented by Zastawny et al. [50], Ouchene et al. [33, 32], Sanjeevi et al. [39, 38], and Chéron et al. [9], Chéron and van Wachem [10]. Zastawny et al. [50] proposed shape-specific correlations for prolate and oblate spheroids, as well as rod-like particles using PR-DNS. They are the most widely used correlations for non-spherical particles in incompressible flows. A detailed analysis of various correlations for prolate spheroids has been performed by Ouchene et al. [33]. They conclude that, while the correlation of Zastawny et al. [50] predict the aerodynamic coefficients for the ellipsoidal particles accurately, the correlations of Zastawny et al. [50] only consider ellipsoidal particles with two different aspect ratios. To address this, Ouchene et al. [32] propose new correlations for drag, lift and pitching torque, which are valid for aspect ratios ranging from $1-32$ of prolate particles. Another recent study on non-spherical particles has been conducted by Sanjeevi et al. [39] using the Lattice-Boltzmann method. The shape-specific correlations proposed by them are applicable for particle Reynolds numbers $\mathrm{Re_{p}}$ ranging from the Stokes limit to 2000, and depend on the particle Reynolds number and the angle of attack of the particle. They extended the correlation to include prolate particles with different aspect ratios, the upper limit of the aspect ratio being 32 [38]. More recently, shape-dependent correlations have been extended to accurate correlations to predict the interaction forces between non-spherical particles and locally uniform and non-uniform fluid flow, including the aspect ratio [15, 31, 38, 9].
@@
Bashforth [3] describes experiments with cannonballs, and shows the effect of compressibility on the drag force in the subsonic, transonic and supersonic regimes ($0.3\leq\mathrm{Ma}\leq 2.0$). In this study, they found the drag force to be proportional to the square of the velocity in the subsonic and supersonic regimes and, proportional to the cube of velocity in the transonic regime. This is an important observation in regard to the effect of compressibility on the drag force. Later, Carlson and Hoglund [8] propose a drag correlation for spherical particles in a compressible flow, also including rarefaction effects. This model is applicable for $\mathrm{Ma}\leq 2$ and $10^{-1}\leq\mathrm{Re_{p}}\leq 10^{3}$. Subsequently, Henderson [17] also propose a Reynolds number and Mach number dependent correlation for spherical particles, valid for a larger range of $\mathrm{Re_{p}}\leq 2\times 10^{4}$ and $\mathrm{Ma}\leq 6$ based on existing experimental data. Loth [23] show that $\mathrm{Re_{p}}$$\approx 45$ separates the compression-dominated ($\mathrm{Re_{p}}\gtrapprox 45$) and rarefaction-dominated ($\mathrm{Re_{p}}\lessapprox 45$) regimes. Loth [23] proposes two different drag models for each regime, in between which a so-called nexus condition exists where, the drag coefficient $C_{\mathrm{D}}$ is independent of the particle Mach and Knudsen numbers. Parmar et al. [34] presents a comparative analysis, using the data from Bailey and Starr [2], and shows that the drag coefficient from Henderson’s model deviates at most 16% due to the interpolations and approximations used by Henderson [17], whereas the drag coefficient from the model of Loth [23] deviates by almost 55% at $\mathrm{Ma}\approx 0.9$. In lieu of this, Parmar et al. [34] proposed an improved correlation for spherical particles for the Reynolds number range $10^{2}<\mathrm{Re_{p}}<2\times 10^{5}$ and Mach numbers of $\mathrm{Ma}\leq 1.75$. They divide the correlation into three parts based on the Mach number regime and the effects of compressibility in each of these regimes: the sub-critical regime ($\mathrm{Ma}\lessapprox 0.6$), the intermediate regime ($0.6\lessapprox\mathrm{Ma}\leq 1$), and the supersonic regime ($1<\mathrm{Ma}\leq 1.75$). The sub-critical regime is characterised by weak compressibility effects and the absence of shock waves. This part of the correlation approaches the correlation of Clift and Gauvin [11] in the limit of $\mathrm{Ma}\rightarrow 0$. In the intermediate Mach regime, weak shocks are formed, due to which the drag coefficient increases with increasing Mach number. In the supersonic regime, a bow shock forms, due to which there is a sharp rise in the drag coefficient with respect to the Mach number. Another noteworthy study on the effects of compressibility for an isolated spherical particle is presented in Nagata et al. [29]. They perform an extensive PR-DNS study to quantify drag, lift and torque for Reynolds numbers in the range of $50-300$ and Mach numbers in the range of $0.3-2.0$. They show that increasing the Mach number leads to an increase of the drag coefficient due to the formation of a detached shock wave in the supersonic regime. They also describe the flow behaviour, by analysing the position of the separation point and the location and length of the re-circulation region downstream of the particle. A comparison of the drag coefficients obtained using PR-DNS with the existing drag correlations [8, 11, 17] reveal that there exists a significant variation among the correlations in transonic and supersonic regimes. In order to investigate the effect of the Mach number and the particle Reynolds number on the flow field, Nagata et al. [27] present an experimental study using Schlieren images for $10^{3}\leq\mathrm{Re_{p}}\leq 10^{5}$ and $0.9\leq\mathrm{Ma}\leq 1.6$. They found that the oscillations of the wake and its size decrease with decreasing $\mathrm{Re_{p}}$ for the considered range of Reynolds numbers. This behaviour is opposite to the one observed for $\mathrm{Re_{p}}\leq 10^{3}$. Moreover, the wake structure amplitude decreases with increasing Mach numbers. Using PR-DNS, Nagata et al. [28] examines the change in drag coefficient due to changing $\mathrm{Re_{p}}$ and $\mathrm{Ma}$ based on the location of the separation point on the sphere.
@@ 2 Methodology
where $\rho$ is the density, $u$ is the fluid velocity, $p$ is the pressure, $\tau$ is the shear stress tensor, $h$ is the specific total enthalpy, $T$ is the temperature and $q$ is the conductive heat flux. This set of equations is closed by the ideal gas equation of state, and the constitutive equation for a Newtonian fluid,
@@
where $\mu$ is the fluid viscosity. The reference temperature is set to $T_{0}=300$ K and the reference pressure to $p_{0}=1.0\times 10^{5}$ Pa. Based on the properties of air, the ratio of specific heats is $\gamma=1.4$ and the specific heat capacity at constant pressure is $c_{p}=1008$ J/(kg K). The thermal conductivity and viscosity of the fluid are calculated for various $\mathrm{Ma}$ and $\mathrm{Re_{p}}$ such that the Prandtl number, $\mathrm{Pr}$, of the gas is maintained at $0.71$. Furthermore, the viscosity is assumed to be constant in the domain. The free stream density is set to $\rho_{\infty}=1.1574$ $\mathrm{kg}/\mathrm{m}^{3}$, whereas the free stream fluid velocity is determined based on the prescribed $\mathrm{Ma}$ and $\mathrm{Re_{p}}$ for each simulation.
@@
The fluid flow over a fixed, isolated particle is simulated in a spherical domain with radius $R$, where the particle lies at the centre of the domain. A spherical domain is chosen because the orientation of the particle relative to the flow can be changed without having to reconstruct the mesh. The desired angle of attack, $\alpha$, which is the angle between the inflow velocity vector and the chord line of the particle, is obtained by rotating the domain boundaries. The hemispherical caps of the domain are set to be the inlet and outlet boundaries, as illustrated in Figure 1. For compressible flows, the inlet hemisphere is prescribed with Dirichlet boundary conditions for velocity, pressure, and temperature, whereas the outlet hemisphere is preescribed with Neumann boundary conditions. The particle boundary is simulated as a no-slip wall. The computational domain is discretised using a hexahedral non-uniform body-conforming mesh, with smaller cells near the particle and increasingly larger cells towards the domain boundary. The radius of the domain is $R=50c$, where $c$ is the semi-major axis of the particle. The particle shapes and their aspect ratios ($\beta=a/b$) considered in the current study are: a prolate spheroid ($\beta=5/2$), an oblate spheroid ($\beta=5$), and a rod-like particle ($\beta=5$). The discretised domains for the considered particle shapes are shown in Figure 2.
@@
Figure 3: Convergence behaviour of the drag coefficient for a rod-like particle using coarse and fine base meshes with optional adaptive mesh refinement (AMR). The left plot corresponds to a Reynolds number of 300, a Mach number of 0.3 and an angle of attack of 0°. The right plot corresponds to a Reynolds number of 300, a Mach number of 2 and an angle of attack of 90°.
Figure 4: Example of a simulation with an AMR mesh, showing the mesh refinement after using the criteria for the nondimensional wall distance at the particle surface, the pressure gradient and vorticity.
Figure 3 shows a representative convergence study for the rod-like particle for a Reynolds number of 300, Mach numbers of 0.3 (left) and 2 (right), and angles of attack of 0°(left) and 90°(right). The crosses represent the results for a coarse base mesh with optional dynamic adaptive mesh refinement (AMR). The criteria for mesh refinement and coarsening are based on the pressure gradient, the vorticity, and the nondimensional wall distance ($y^{+}$). The boundary layer is resolved using a maximum $y^{+}\lessapprox 1$, since no wall-treatment is applied and the boundary layer is fully resolved. 1 or 2 levels of AMR refinement mean that the initial coarse cells are subdivided once or twice, resulting in a finer mesh where needed. This hierarchical refinement ensures that the mesh is only highly resolved in areas where it is necessary, optimising computational resources while maintaining accuracy. Since the number of cells for the simulations with AMR varies, the reported number of cells with AMR corresponds to the number of cells when the aerodynamic force and torque coefficients are converged. The circles represent the results for a fine base mesh with optional AMR. Using the AMR significantly reduces the computational cost by only adapting the cells close to the particle, so that the wake behind the particle and any discontinuities such as shock waves are captured. Therefore, the coarse base mesh with two levels of AMR refinement is used for all further simulations, since the error is below 5 percent compared to the fine base mesh with AMR, but with considerably lower computational costs. An exemple of a simulation with an AMR mesh is shown in Figure 4. The coarse base mesh with 2 levels of AMR for each of the shapes contains a maximum of approximately $2\times 10^{5}$ to $3\times 10^{5}$ hexahedral cells.
The forces acting on the surface of the particles and their corresponding coefficients are computed by
${F_{i}=\int_{S}\left(-p\delta_{ij}+\mu\left[\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right]-\frac{2}{3}\mu\frac{\partial u_{k}}{\partial x_{k}}\delta_{ij}\right)n_{j}\mathrm{d}S},$ (5)
${C_{\mathrm{D}}=\frac{F_{\mathrm{D}}}{\frac{1}{2}\rho_{\infty}u_{\infty}^{2}A_{\mathrm{p}}},\,\,C_{\mathrm{L}}=\frac{F_{\mathrm{L}}}{\frac{1}{2}\rho_{\infty}u_{\infty}^{2}A_{\mathrm{p}}}},$ (6)
where $S$ is the surface of the particle, $\mu$ is the viscosity of the fluid, $\vec{n}$ is the normal vector to the particle surface, $\rho_{\infty}$ is the free stream density, $u_{\infty}$ is the free stream velocity, $F_{\mathrm{D}}$ and $F_{\mathrm{L}}$ are the forces acting parallel and normal to the particle surface, respectively, and $A_{\mathrm{p}}=\pi d_{\mathrm{p}}^{2}/4$ is the reference area of the particle based on the equivalent diameter of a sphere of the same volume, $d_{\mathrm{p}}$. The torque exerted on an axi-symmetric particle can be decomposed into two principal components: the pitching or hydrodynamic torque, and the rotational torque. The pitching torque originates from the displacement between the particle centre of pressure within the fluid and its centre of mass. This displacement results from the varying distance between the centre of pressure’s position and the centre of mass. The centre of mass remains fixed in the reference frame of a rigid particle. Under a uniform flow configuration, the displacement of the centre of pressure is noted when the angle of attack between the particle and the inlet flow direction ranges from $0\degree$ to $90\degree$. Specifically, at angles of attack of $0\degree$ and $90\degree$, the pressure distribution over the particle surface is symmetric, leading to the coincidence of the centre of pressure with the centre of mass, thereby eliminating the pitching torque. The pitching torque coefficient is computed based on the drag coefficient as [36]
$C_{\mathrm{M}}=C_{\mathrm{D}}\cdot\sin{\alpha}\cdot\cos{\alpha},$ (7)
with $\alpha$ being the angle of attack.
@@
The time step is constrained by setting the maximum Courant-Friedrichs-Lewy (CFL) number to 0.3. The simulations are run until a steady state is observed in the temporal evolution of the aerodynamic force coefficients. The residuals have a convergence criterion of $O(10^{-8})$ and a total conservation error of $O(10^{-3})$ is achieved. All aerodynamic force coefficients shown in the results section are mean values of the last 1000 time steps.

### v2

@@ Abstract
This study presents particle-resolved direct numerical simulations using three-dimensional body-fitted hexahedral meshes to investigate the aerodynamic force and torque coefficients of non-spherical particles in compressible flows. The simulations focus on three particle shapes: a prolate spheroid, an oblate spheroid, and a rod-like particle, across a range of Mach numbers (0.3 to 2.0), angles of attack (0°to 90°), and particle Reynolds numbers (100 to 300). Results show that the particle shape significantly impacts the aerodynamic forces on a particle in a compressible flow, with oblate spheroids exhibiting the highest drag, lift, and torque values. Correlations for these aerodynamic coefficients of the particles in a compressible flow are developed and validated. These correlations advance multiphase flow modeling by improving the accuracy of point-particle simulations for non-spherical particles in compressible flows.
@@ 1 Introduction
Particle-resolved direct numerical simulations (PR-DNS) of particle-laden flows require tremendous computational resources. However, a point-particle approach can significantly reduce the computational power required to model particle-laden flows. This approach requires force models to determine the forces exerted on the particles by the surrounding fluid, as well as by contact with other particles [26]. In this context, several force models have been developed which are applicable for various flow regimes and volume fractions [44, 50, 23, 24, 39, 38, 11, 17, 34, 9]. In order to derive these correlations, it is essential to first understand the behaviour of an isolated particle and the forces experienced by the particle in various flow regimes. These force correlations can then be extended to multi-particle systems.
Forces acting on an isolated particle depend on several factors, such as the particle Reynolds number, the Mach number, the fluid properties, the particle shape, and angle of attack between the mean fluid flow and the axis of the particle. Predicting the forces that act on a particle in a fluid flow and predicting its behaviour have been the subject of studies for decades [43, 48, 17, 36, 19, 25, 44, 29, 30, 27, 47]. One of the first analytical models to predict the drag of a spherical particle for particle Reynolds numbers in the Stokes regime ($\mathrm{Re}_{\mathrm{p}}\ll 1$) was given by Stokes [43] as $C_{\mathrm{D}}=24/\mathrm{Re_{p}}$ where $C_{\mathrm{D}}$ is the drag coefficient. This model is the basis for later proposed drag models [22, 8, 16, 37], applicable to various flow regimes and practical applications.
@@
Since particles are not limited to spherical shapes, the proposed models also included non-spherical particle shapes. Some of the most notable correlations for non-spherical particle shapes were those presented by Zastawny et al. [50], Ouchene et al. [33, 32], Sanjeevi et al. [39, 38], and Chéron et al. [9], Chéron and van Wachem [10]. Zastawny et al. [50] proposed shape-specific correlations for prolate and oblate spheroids, as well as rod-like particles using PR-DNS. They are the most widely used correlations for non-spherical particles in incompressible flows. A detailed analysis of various correlations for prolate spheroids has been performed by Ouchene et al. [33]. They conclude that, while the correlation of Zastawny et al. [50] predict the aerodynamic coefficients for the ellipsoidal particles accurately, the correlations of Zastawny et al. [50] only consider ellipsoidal particles with two different aspect ratios. To address this, Ouchene et al. [32] propose new correlations for drag, lift and torque, which are valid for aspect ratios ranging from $1-32$ of prolate particles. Another recent study on non-spherical particles has been conducted by Sanjeevi et al. [39] using the Lattice-Boltzmann method. The shape-specific correlations proposed by them are applicable for particle Reynolds numbers $\mathrm{Re_{p}}$ ranging from the Stokes limit to 2000, and depend on the particle Reynolds number and the angle of attack of the particle. They extended the correlation to include prolate particles with different aspect ratios, the upper limit of the aspect ratio being 32 [38]. More recently, shape-dependent correlations have been extended to accurate correlations to predict the interaction forces between non-spherical particles and locally uniform and non-uniform fluid flow, including the aspect ratio [15, 31, 38, 9].
@@
Bashforth [3] describes experiments with cannonballs, and shows the effect of compressibility on the drag force in the subsonic, transonic and supersonic regimes ($0.3\leq\mathrm{Ma}\leq 2.0$). In this study, they found the drag force to be proportional to the square of the velocity in the subsonic and supersonic regimes and, proportional to the cube of velocity in the transonic regime. This is an important observation in regard to the effect of compressibility on the drag force. Later, Carlson and Hoglund [8] propose a drag correlation for spherical particles in a compressible flow, also including rarefaction effects. This model is applicable for $\mathrm{Ma}\leq 2$ and $10^{-1}\leq\mathrm{Re_{p}}\leq 10^{3}$. Subsequently, Henderson [17] also propose a Reynolds number and Mach number dependent correlation for spherical particles, valid for a larger range of $\mathrm{Re_{p}}\leq 2\times 10^{4}$ and $\mathrm{Ma}\leq 6$ based on existing experimental data. Loth [23] show that $\mathrm{Re_{p}}$$\approx 45$ separates the compression-dominated ($\mathrm{Re_{p}}\gtrapprox 45$) and rarefaction-dominated ($\mathrm{Re_{p}}\lessapprox 45$) regimes.
@@
Another noteworthy study on the effects of compressibility for an isolated spherical particle is presented in Nagata et al. [29]. They perform an extensive PR-DNS study to quantify drag, lift and torque for Reynolds numbers in the range of $50-300$ and Mach numbers in the range of $0.3-2.0$. They show that increasing the Mach number leads to an increase of the drag coefficient due to the formation of a detached shock wave in the supersonic regime. A comparison of the drag coefficients obtained using PR-DNS with the existing drag correlations [8, 11, 17] reveal that there exists a significant variation among the correlations in transonic and supersonic regimes. In order to investigate the effect of the Mach number and the particle Reynolds number on the flow field, Nagata et al. [27] present an experimental study using Schlieren images for $10^{3}\leq\mathrm{Re_{p}}\leq 10^{5}$ and $0.9\leq\mathrm{Ma}\leq 1.6$. They found that the oscillations of the wake and its size decrease with decreasing $\mathrm{Re_{p}}$ for the considered range of Reynolds numbers. This behaviour is opposite to the one observed for $\mathrm{Re_{p}}\leq 10^{3}$. Moreover, the wake structure amplitude decreases with increasing Mach numbers. Using PR-DNS, Nagata et al. [28] examines the change in drag coefficient due to changing $\mathrm{Re_{p}}$ and $\mathrm{Ma}$ based on the location of the separation point on the sphere.
@@ 2 Methodology
where $\rho$ is the density, $\mathbf{u}$ is the fluid velocity, $p$ is the pressure, $\boldsymbol{\tau}$ is the shear stress tensor, $h$ is the specific total enthalpy, and $\mathbf{q}$ is the conductive heat flux. This set of equations is closed by the ideal gas equation of state, and the constitutive equation for a Newtonian fluid,
@@
where $\mu$ is the fluid viscosity. The reference temperature is set to $\theta_{0}=300$ K and the reference pressure to $p_{0}=1.0\times 10^{5}$ Pa. Based on the properties of air, the ratio of specific heats is $\gamma=1.4$ and the specific heat capacity at constant pressure is $c_{p}=1008$ J/(kg K). The thermal conductivity and viscosity of the fluid are calculated for various $\mathrm{Ma}$ and $\mathrm{Re_{p}}$ such that the Prandtl number, $\mathrm{Pr}$, of the gas is maintained at $0.71$. Furthermore, the viscosity is assumed to be constant in the domain. The free stream density is set to $\rho_{\infty}=1.1574$ $\mathrm{kg}/\mathrm{m}^{3}$, whereas the free stream fluid velocity is determined based on the prescribed $\mathrm{Ma}$ and $\mathrm{Re_{p}}$ for each simulation.
@@
The fluid flow over a fixed, isolated particle is simulated in a spherical domain with radius $R$, where the particle lies at the centre of the domain. A spherical domain is chosen because the orientation of the particle relative to the flow can be changed without having to reconstruct the mesh. The desired angle of attack, $\alpha$, which is the angle between the inflow velocity vector and the chord line of the particle, is obtained by rotating the domain boundaries. The hemispherical caps of the domain are set to be the inlet and outlet boundaries, as illustrated in Figure 1. For compressible flows, the inlet hemisphere is prescribed with Dirichlet boundary conditions for velocity, pressure, and temperature, whereas the outlet hemisphere is preescribed with Neumann boundary conditions. The particle boundary is simulated as a no-slip wall with an adiabatic boundary condition for temperature. The computational domain is discretised using a hexahedral non-uniform body-conforming mesh, with smaller cells near the particle and increasingly larger cells towards the domain boundary. The radius of the domain is $R=50c$, where $c$ is the semi-major axis of the particle. The particle shapes and their aspect ratios ($\beta=a/b$) considered in the current study are: a prolate spheroid ($\beta=5/2$), an oblate spheroid ($\beta=5$), and a rod-like particle ($\beta=5$). The discretised domains for the considered particle shapes are shown in Figure 2.
@@
In A, Figures 15 to 17 show convergence studies for the three considered non-spherical particle shapes for a Reynolds number of 300, Mach numbers of 0.3 and 2, and an angle of attack of 0°. Based on the results of the convergence studies in A, the medium-sized mesh is chosen for all simulations of all three non-spherical shapes as it gives the best trade-off between accuracy and computational cost. Compared to the fine mesh, the relative error of the drag coefficient for the medium is below 5% for all convergence studies, except for the prolate spheroid at $\mathrm{Ma}=2$ where the error is 6%. The ratio of the equivalent diameter of a sphere of the same volume to the smallest mesh spacing at the particle surface is 75 ($\Delta x=0.072\,\mathrm{m}$) for the prolate spheroid, 115 ($\Delta x=0.051\,\mathrm{m}$) for the oblate spheroid and 50 ($\Delta x=0.0019\,\mathrm{m}$) for the rod-like particle, which leads to a resolved boundary layer for all 3 non-spherical particle shapes.
The force $\mathbf{F}$ and the torque $\mathbf{T}$ acting on the surface of the particles are computed by
${F_{i}=\int_{S}\tau_{ij}n_{j}\mathrm{d}S},$ (5)
$T_{i}=\int_{S}\epsilon_{ijk}r_{j}\tau_{km}n_{m}\,\mathrm{d}S,$ (6)
where $S$ is the surface of the particle, $\mathbf{n}$ is the normal vector to the particle surface, $\mathbf{r}$ is the vector from the particle center to the point where the force is applied on the particle surface, and $\epsilon$ is Levi-Civita symbol for the cross product. The corresponding coefficients are defined as
$C_{\mathrm{D}}=\frac{F_{\mathrm{D}}}{\frac{1}{2}\rho_{\infty}u_{\infty}^{2}A_{\mathrm{p}}},\,\,C_{\mathrm{L}}=\frac{F_{\mathrm{L}}}{\frac{1}{2}\rho_{\infty}u_{\infty}^{2}A_{\mathrm{p}}},\,\,C_{\mathrm{M}}=\frac{T_{\mathrm{p}}}{\frac{1}{2}\rho_{\infty}u_{\infty}^{2}\pi d_{\mathrm{p}}^{3}/8},$ (7)
where $F_{\mathrm{D}}$ and $F_{\mathrm{L}}$ are the forces acting parallel and normal to the flow direction, respectively, $T_{\mathrm{p}}$ is the pitching torque acting on the particle, $\rho_{\infty}$ is the free stream density, $u_{\infty}$ is the free stream velocity, and $A_{\mathrm{p}}=\pi d_{\mathrm{p}}^{2}/4$ is the reference area of the particle based on the equivalent diameter of a sphere of the same volume, $d_{\mathrm{p}}$.
@@
The time step is constrained by setting the maximum Courant-Friedrichs-Lewy (CFL) number to 0.3. The simulations are run until a steady state is observed in the temporal evolution of the aerodynamic force coefficients. The residuals have a convergence criterion of $O(10^{-8})$ and a total conservation error of $O(10^{-3})$ is achieved. All aerodynamic force coefficients shown in the results section are mean values of at least the last 1000 time steps.